# CAVE: Controllable Authorship Verification Explanations

### v2

@@ Appendix B PromptAV and CoT
Tables B and B show the prompts used to apply CoT and PromptAV respectively. The prompts are taken from Hung et al. (2023); further note that the prompt is uniform across datasets.
@@ 8. Other Aspects:
PromptAV (and by extension, CoT) which serve as the inspiration for Cave produce outputs which cannot be automatically checked for consistency; these prompts produces a free-text explanation followed by a final score which can be thresholded to produce a YES/NO Av label. As observed in the experiments of Hung et al. (2023), models tend to produce low confidence scores with PromptAV, which means that for better accuracy, the optimal threshold in their experiments with GPT-3.5-Turbo was found to be 0.2-0.3, which will most probably not be consistent with the free text rationale (that is, if the rationale says that it thinks it is not confident that the authors are the same, and the confidence is 0.25, the automatic label prediction will still assume that the predicted label is YES). Further, with just the score as the label, it is still possible that we misconstrue the label because the confidence score and the model’s predicted label in the text are different, as shown in the examples in Tables 8 and 9.
Appendix C Hyperparameters
@@
We download the pre-trained LLaMa-3-8B from HuggingFace (https://huggingface.co/meta-llama/Meta-Llama-3-8B) and train it for all our experiments. We train it with Low-Rank Adaption (LoRA; Hu et al., 2021), a parameter-efficient training method which offers lightweight training no additional inference latency, unlike other existing parameter-efficient methods Houlsby et al. (2019). For single-dataset Cave, we train LLaMa-3-8B for $10$ epochs; for multi-dataset Cave, we train for $5$ epochs. We use no validation set, and instead use the last checkpoint obtained from training; since we have no gold-standard explanations, we opt against using the GPT-4-Turbo generated data as the validation set. We experiment with different values of ‘r’ and ‘alpha’ for LoRA (refer to https://huggingface.co/docs/peft/main/en/conceptual_guides/lora for their definitions). For IMDb62 and FanFiction, we use r=128 and alpha=256; for Blog-Auth and for the multi-dataset case, we use r=64 and alpha=128. Note that all our training experiments for LLaMa-3-8B take $\sim$12 hours, and all our test set inferences take $\sim$7-8 hours on a single NVIDIA Quadro RTX 8000 GPU. We report single-run results for all our experiments.
@@
punctuation style 0 1 15 9 15 19 - - -
special characters/ 9 11 6 14 13 17 12 13 9
@@
acronyms/abbr. 14 11 4 15 14 17 18 15 3
writing style 18 18 19 20 19 20 14 17 17
expressions/idioms 16 15 13 18 16 18 17 17 11
tone and mood 19 17 19 19 20 20 17 17 18
sentence structure 18 18 14 15 19 20 16 16 10
any other 18 17 16 18 20 13 17 19 12
@@
Appendix D Cave with different offline models
@@
LLaMa-3-8B Single-dataset 75.7* 0.97 60.4 0.94 60.2 0.81
Multi-dataset 74.9 0.99* 57.9 0.96 58.6 0.91*
Mistral-7B Single-dataset 66.0 0.91 58.0 0.87 53.6 0.96
Multi-dataset 74.3 0.96 59.6 0.92 57.4 0.83
Stable-LM-Zephyr-3B Single-dataset 42.5 0.67 47.7 0.81 48.3 0.66
Multi-dataset 34.9 0.47 32.7 0.51 17.5 0.22
Gemma-2-2B Single-dataset 24.5 0.36 32.3 0.55 27.9 0.42
Multi-dataset 53.9 0.69 36.9 0.58 52.2 0.77
@@
Table 13: Empirical Results comparing Cave trained with local models of different sizes. Mistral-7B is the next best performing model after LLaMa-3-8B. Gemma-2-2B and Stable-LM-Zephyr-3B however have lower performances; further, the former’s performance increases with multi-dataset training while the latter’s decreases.
@@
We report experiments with local models of varying sizes in Table 13. After LLaMa-3-8B, Mistral-7B is the next best performing model. The smaller models Gemma-2-2B and Stable-LM-Zephyr-3B however have lower performances; further, the former’s performance increases with multi-dataset training while the latter’s decreases.
Appendix E Uncertainty of Cave
Our metric Consistency serves the dual purpose of measuring consistency as well as uncertainty of the model in its prediction. We mark that a rationale’s $\textsc{Consistency}=1$ if and only if the aggregated intermediate labels, final score and final label all match. If they don’t all match, we deem the output as inconsistent and hence uncertain. Hence, the consistency numbers in Table 3 reflect the uncertainty of the model in its output. Specifically, we note that Cave’s outputs show high consistencies (the highlighted/last two rows in Table 3) indicating that Cave is highly certain of its outputs.
Appendix F Human Evaluation
@@
Table B shows the instructions given to the annotators regarding the properties to be analyzed and Table 11 shows a sample human evaluation. We perform an in-house human evaluation, and annotators used were proficient in English and well aware of the high complexity of the task.
@@
Table 12 shows the human pilot study numbers for individual datasets. The aggregate of these numbers were used for the overall analysis in Section 3.5.
Appendix G Random Errors due to Computational Limitations
@@
We acknowledge random errors (i.e., not systematic) that could have occurred in our results owing to limitations of language models and computational resources. Our method uses a language model to read two input documents and provide an output; however, in practice, language models have a fixed maximum length of input due to model limitations as well as GPU memory limitations. This means that if the two input documents together (along with the instructions) do not fit into the input, it becomes necessary to truncate them, which leads to loss of information of their latter parts. This issue is exacerbated when we move towards using in-house models on user servers (which typically have a shorter context window than API-based models like GPT-4-Turbo). In our experiments with LLaMa-3-8B, we used just 1 GPU (to have a realistic user environment), which means that while training we had to restrict the input to be of maximum length 768768768768121212note: during inference, we did not limit the size of the input, but this could pose a problem if the documents became much longer, which could have affected the training for the longer documents. Further, even for generating the training data from GPT-4-Turbo for FanFiction, the excessive size of the documents made the GPT-4-Turbo generations economically infeasible; we cut our the input to to a maximum of 300 words per document, which could have created errors/incomplete rationales that eventually propagated to our downstream training of LLaMa-3-8B.
Appendix H Out-of-domain inference
@@
IMDb62 57.7 0.95 58.7 0.92
Blog-Auth 73.8 0.93 50.6 0.74
FanFiction 68.6 0.70 57.4 0.83
Table 14: Out of domain performance by single-dataset Cave models
Finally, we evaluate the single dataset Cave models on the other datasets’ test sets (refer Table 14); we notice robust out-of-domain performance. Notably, the IMDb62 Cave model performs extremely well across the board, with competitive accuracies and consistencies for all datasets. This is a significant result, as it shows the generalizability of this approach to unseen Av datasets: although we use fine-tuning to improve performance over a zero-shot approach, it is not fine-tuning specifically to the domain that is most critical but rather fine-tuning to the structured format of Cave’s output.

### v3

@@ Appendix C PromptAV and CoT
Tables C and C show the prompts used to apply CoT and PromptAV respectively. The prompts are taken from Hung et al. (2023); further note that the prompt is uniform across datasets.
@@ 8. Other Aspects:
PromptAV (and by extension, CoT) which serve as the inspiration for Cave produce outputs which cannot be automatically checked for consistency; these prompts produces a free-text explanation followed by a final score which can be thresholded to produce a YES/NO Av label. As observed in the experiments of Hung et al. (2023), models tend to produce low confidence scores with PromptAV, which means that for better accuracy, the optimal threshold in their experiments with GPT-3.5-Turbo was found to be 0.2-0.3, which will most probably not be consistent with the free text rationale (that is, if the rationale says that it thinks it is not confident that the authors are the same, and the confidence is 0.25, the automatic label prediction will still assume that the predicted label is YES). Further, with just the score as the label, it is still possible that we misconstrue the label because the confidence score and the model’s predicted label in the text are different, as shown in the examples in Tables 13 and 14.
Appendix D GPT-4-Turbo silver data analysis
In Table 15 we show train dataset sizes at every stage of data generation and preprocessing.
Further, we also perform manual analysis of GPT-4-Turbo’s silver rationales. For the most part, GPT-4-Turbo generates high-quality rationales; the most frequent issue observed is hallucination of details, especially punctuation marks. Out of the 60 samples analyzed (20 from each dataset), 10 samples showed hallucination in their rationale for punctuation style, 16 samples showed factual incorrectness across varying linguistic features, and 6 samples showed inconsistency of rationale with intermediate label in one linguistic feature (varying across samples) each.
Dataset Initial size Filter for Structure Filter for float score Filter for Cons-R-L
IMDb62 3400 3396 3255 2354
Blog-Auth 3600 3600 3146 2724
FanFiction 3600 3383 2887 2159
Table 15: Train dataset sizes at every stage of data generation and preprocessing. The ‘Initial size’ column shows the number of datapoints generated from GPT-4-Turbo, and each succeeding column shows the number of datapoints remaining after the different filtering methods (in order of application).
Appendix E Hyperparameters
@@
We download the pre-trained LLaMa-3-8B from HuggingFace (https://huggingface.co/meta-llama/Meta-Llama-3-8B) and train it for all our experiments. We train it with Low-Rank Adaption (LoRA; Hu et al., 2021), a parameter-efficient training method which offers lightweight training no additional inference latency, unlike other existing parameter-efficient methods Houlsby et al. (2019). We train LLaMa-3-8B for $10$ epochs. We use no validation set, and instead use the last checkpoint obtained from training; since we have no gold-standard explanations, we opt against using the GPT-4-Turbo generated data as the validation set. We experiment with different values of ‘r’ and ‘alpha’ for LoRA (refer to https://huggingface.co/docs/peft/main/en/conceptual_guides/lora for their definitions). We use r=128 and alpha=256. Note that all our training experiments for LLaMa-3-8B take $\sim$12 hours, and all our test set inferences take $\sim$7-8 hours on a single NVIDIA Quadro RTX 8000 GPU. We report single-run results for all our experiments.
Why not LLaMa-3-8B-Instruct for fine-tuning? As explained in the documentation for LLaMa-3-8B-Instruct in huggingface https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, “Instruction tuned models are intended for assistant-like chat, whereas pretrained models can be adapted for a variety of natural language generation tasks”. Hence, we provide zero-shot and few-shot baselines with LLaMa-3-8B-Instruct whereas we provide fine-tuned results on LLaMa-3-8B.
Document truncation. When we truncate documents for GPU memory considerations in the training of Cave, we always ensure that equal amounts of both documents are present in the input. Further, we always use the first chunk of the document as input and not a random chunk.
@@
punctuation style 15 31 43 25 38 45 23 34 37
special characters/ 31 40 47 23 37 46 33 39 43
@@
acronyms/abbr. 38 41 47 25 37 45 47 46 43
writing style 47 45 46 44 48 49 45 44 50
expressions/idioms 44 48 49 40 43 45 46 46 44
tone and mood 47 48 48 46 47 48 47 43 49
sentence structure 45 48 46 42 45 48 39 40 47
any other 48 43 38 46 38 37 48 46 45
@@
Appendix F Cave with different offline models, Practical advantages of Cave
@@
LLaMa-3-8B Fine-tune 74.1 0.99 60.6 1.00 62.6 0.99
Mistral-7B Fine-tune 73.7 0.99 59.2 0.99 57.7 0.98
Stable-LM-Zephyr-3B Fine-tune 50.8 0.71 47.7 0.81 42.0 0.68
@@
Table 19: Empirical Results comparing Cave trained with local models of different sizes. Mistral-7B is the next best performing model after LLaMa-3-8B. Stable-LM-Zephyr-3B however has a lower performance.
@@
We report experiments with local models of varying sizes in Table 19. After LLaMa-3-8B, Mistral-7B is the next best performing model.
Practical Advantages:
Cave also demonstrates practical advantages over GPT-4-Turbo such as (1) Energy: LLaMa-3-8B is just 8B parameters, whereas GPT-4-Turbo is orders of magnitude bigger - Cave hence consumes much less energy per datapoint, and hence is more affordable and efficient, (2) Financial cost: Cave is also more financially affordable since it can be hosted on a local server; GPT-4-Turbo or any API based language model would cost much more since it charges for every input/output token, (3) Speed: On average, Cave takes 3 minutes per datapoint for inference, whereas GPT-4-Turbo takes 4.5 minutes.
F.1 Human Pilot
Appendix G Human Evaluation
@@
Table C shows the instructions given to the annotators regarding the properties to be analyzed and Table 17 shows a sample human evaluation. We perform an in-house human evaluation, and annotators used were proficient in English and well aware of the high complexity of the task.
@@
Table 18 shows the human pilot study numbers for individual datasets. The aggregate of these numbers were used for the overall analysis in Appendix F.1.
Appendix H Random Errors due to Computational Limitations
@@
We acknowledge random errors (i.e., not systematic) that could have occurred in our results owing to limitations of language models and computational resources. Our method uses a language model to read two input documents and provide an output; however, in practice, language models have a fixed maximum length of input due to model limitations as well as GPU memory limitations. This means that if the two input documents together (along with the instructions) do not fit into the input, it becomes necessary to truncate them, which leads to loss of information of their latter parts. This issue is exacerbated when we move towards using in-house models on user servers (which typically have a shorter context window than API-based models like GPT-4-Turbo). In our experiments with LLaMa-3-8B, we used just 2 GPUs (to have a realistic user environment), which means that while training we had to restrict the input to be of maximum length 1280128012801280161616note: during inference, we did not limit the size of the input, but this could pose a problem if the documents became much longer, which could have affected the training for the longer documents. Further, even for generating the training data from GPT-4-Turbo for FanFiction, the excessive size of the documents made the GPT-4-Turbo generations economically infeasible; we cut our the input to to a maximum of 300 words per document, which could have created errors/incomplete rationales that eventually propagated to our downstream training of LLaMa-3-8B.
Appendix I Out-of-domain inference
@@
IMDb62 59.9 1.0 61.2 0.99
Blog-Auth 71.3 0.96 45.3 0.77
FanFiction 68.8 1.0 59.3 1.0
Table 20: Out of domain performance by Cave
Finally, we evaluate Cave models on the other datasets’ test sets (refer Table 20); we notice robust out-of-domain performance especially by the IMDb62 and FanFiction models. This is a significant result, as it shows the generalizability of this approach to unseen Av datasets: although we use fine-tuning to improve performance over a zero-shot approach, it is not fine-tuning specifically to the domain that is most critical but rather fine-tuning to the structured format of Cave’s output.
Appendix J Post-hoc Interpretability
We finally clarify that Cave was designed to serve post-hoc rationalization, and not mechanistic (i.e., low-level mathematical) understanding of the language model’s workings. We quote from Lipton (2018) to emphasize that Cave’s improved post-hoc explainability leads to improved utility to (1) downstream users, and (2) researchers, who can leverage Cave’s accessible outputs for detailed analyses, to address errors and develop improved models: “while post-hoc explanations often do not elucidate precisely how a model works, they may nonetheless confer useful information for practitioners and end users of machine learning”.
Appendix K More rationale properties
In addition to consistency, there are several properties necessary for high quality rationales Joshi et al. (2023); Ramnath et al. (2023). Some of these are easily verifiable by humans, even if there exist no prevalent automatic methods to measure them: for eg:- factual correctness, hallucination in the rationale, etc. that we measured in our human evaluation (§F.1). There are also properties that are hard to verify even by human evaluators: for eg:- completeness of the rationale, i.e., has the rationale missed any significant (dis)similarity between the documents?. In fact, completeness does not have a comprehensive definition even for simpler tasks like multi-choice QA Ramnath et al. (2023). In the future, we would like to work on automatic metrics and model training for these properties specifically for Av.
Appendix L Analyzing cases where PromptAV succeeds but Cave fails
We consider the $88$ samples of the Blog-Auth test set where PromptAV succeeds but Cave fails. From qualitative analysis of a few samples, we find out the following (non-exhaustive) patterns:
  • Cases where the gold label is ‘YES’, and both methods fail: Cave wrongly predicts a label of ‘NO’ with a final score of $<0.5$. PromptAV gives the wrong reasoning, but concludes by predicting a confidence score $>0.5$ that the authors are different which leads to a misinterpretation that PromptAV was right. We provide an example of this error by PromptAV in Appendix C.
  • Differences in perceiving patterns such as idioms, acronyms: Both methods generate similar explanations for abstract features such as writing style, tone & mood. But when it comes to token-level features such as acronyms, idioms, etc., the methods sometimes predict different subsets, which leads to differing final predictions. This comes under the issue of completeness of rationale, which we discuss in the Limitations section and in Appendix K. In this particulat situation, both models have provided different incomplete rationales, which eventually leads to different labels.